the sum of measures of the two acute angle is 110. the supplement of the smaller angle is fifteen times the complement of the larger angle. find the measures of the acute angle .
Answer by mananth(16946)

let one acute angle be x deg.
the other will be 110-x deg.
..
180-x = 15(90-(110-x))
180-x = 15(90-110+x)
180-x=15(-20+x)
180-x=-300+15x
-15x-x=-300-180
-16x=-480
x=30 deg.